# Three-term Machin-type formulae

### v5

@@ 1. Introduction
    If $m_{1}m_{2}\geq 10^{10}$ and $m_{2}$ does not divide $x_{i}^{2}+1$ for some $i$, then $m_{1}<2.549\times 10^{24}$, $m_{2}<\exp(295490)$, $x_{i}<\exp(5.07\times 10^{12})$, and $\left\lvert y_{i}\right\rvert<1.316\times 10^{19}$. Moreover, $m_{2}<e^{40000}$ if additionally $m_{1}\geq 9.203\times 10^{22}$.
@@ 3. A lower bound for linear forms of three logarithms
(92) $\begin{split}{b^{\prime}}=&\frac{b_{1}}{9.05\pi+2D\operatorname*{h}(\alpha)}+\frac{b_{2}}{9.05\pi},\\ D=&[{\mathbb{Q}}(\alpha):{\mathbb{Q}}]/2,\\ a=&9.05\pi+2D\operatorname*{h}(\alpha),\\ H=&\max\{17,D,D(\log{b^{\prime}}+2.96)+0.01\}.\end{split}$
@@
(93) $\log\left\lvert\Lambda_{1}\right\rvert>-2.7717aH^{2}.$
@@
This follows from Theorem 1 of [6] with $D=1$ by setting $\mu=0.59$, $\rho_{1}=18.1$, $\alpha_{1}=i$, $\alpha_{2}=\alpha$, $a_{1}=\rho\pi/2$, $a_{2}=2D\operatorname*{h}(\alpha)+(\rho\pi/2)$, $L=\left\lfloor H+\sqrt{H^{2}+(1/4)}+(1/2)\right\rfloor$, $k=(v_{1}(L)+\sqrt{v_{1}(L)^{2}+4v_{0}(L)v_{2}(L)})^{2}/(2v_{2}(L))^{2}$ to be the positive real number such that $\sqrt{k}$ satisfies the quadratic equation $v_{2}(L)k-v_{1}(L)\sqrt{k}-v_{0}(L)=0$, where
@@
Then, we can confirm that $0.27470.2432<\sqrt{k}<$ 0.2747, $H>7.5$, $15\leq L<0.92h$, and $K>700$ and then the condition (2) of Theorem 1 of [6]. Moreover, we can also confirm that $R_{2}-1<2a_{2}H^{2}<b_{1}$ and $S_{2}-1<2a_{1}H^{2}<b_{2}$. Hence the condition (1) of Theorem 1 of [6] also holds. Then, confirming that $\sqrt{k}L<0.23865H$, Theorem 1 of [6] yields that $\log\left\lvert\Lambda_{1}\max\{LSe^{LS\left\lvert\Lambda_{1}\right\rvert/(2b_{2})}/(2b_{2}),LRe^{LR\left\lvert\Lambda_{1}\right\rvert/(2b_{1})}/(2b_{1})\}\right\rvert>-2.7706aH^{2}$ and then $\log\left\lvert\Lambda_{1}\right\rvert>-2.7717aH^{2}$. For the detail of the proof, see [22]. ∎
@@ 5. Proof of the Theorem
We begin by showing that $m_{1}<M_{1}$ with $M_{1}=2.549\times 10^{24}$. Assume that $m_{2}\geq m_{1}\geq M_{1}$. Now, Theorem 4.1 allows us to take $A_{4}=1752.029$, $A_{5}=10545512.8\cdots$, and $\rho=7.292$.
@@
Now, if $m_{1}\leq 6.5\times 10^{8}G^{2}$, then, (120) gives $m_{1}<6.5\times 10^{8}A_{4}^{2}(\log^{2}m_{1})\log^{4}(10545513\log m_{1})$ and therefore we must have $m_{1}<2.549\times 10^{24}<M_{1}$. Thus, we may assume that $m_{2}>m_{1}>6.5\times 10^{8}G^{2}$.
@@
(123) $\left\lvert\frac{r\pi}{4}-y_{1}\arctan\frac{1}{x_{1}}\right\rvert<\frac{k_{1}\ell_{3}}{x_{2}-1}+\frac{k_{1}\ell_{2}}{x_{3}-1}<\frac{(2+10^{-5})G}{\sqrt{m_{2}}}<7.9\times 10^{-5}.$
@@
(124) $\frac{\left\lvert y_{1}\right\rvert}{\sqrt{m_{1}}}>(1-10^{-9})\left(\frac{\pi}{4}-7.9\times 10^{-5}\right)>0.785319.$
@@
(129) $\log\left\lvert\Lambda_{1}\right\rvert>-2.7717aH^{2},$
@@
(130) $\begin{split}a=&9.05\pi+\log m_{1}\\ H=&\max\left\{17,2.97+\log\left(\frac{\left\lvert r\right\rvert}{9.05\pi+\log m_{1}}+\frac{\left\lvert y_{1}\right\rvert}{9.05\pi}\right)\right\}.\end{split}$
@@
(131) $\left\lvert\frac{r\pi}{4}\right\rvert-7.9\times 10^{-5}\leq\frac{\left\lvert y_{1}\right\rvert}{x_{1}-1}<\frac{A_{4}(\alpha\log m_{2})\log^{2}(10545513\log m_{1})}{m_{1}^{1/2}}.$
@@
(132) $\frac{\left\lvert r\right\rvert}{9.05\pi+\log m_{1}}+\frac{\left\lvert y_{1}\right\rvert}{9.05\pi}<A_{4}\psi_{3}(m_{1})(\alpha\log m_{2})\log^{2}(10545513\log m_{1}),$
@@
(133) $\begin{split}\psi_{3}(m_{1})=&\leavevmode\nobreak\ \frac{4}{(9.05\pi+\log m_{1})\pi m_{1}^{1/2}}+\frac{1}{9.05\pi}\\ &\leavevmode\nobreak\ +\frac{3.16\times 10^{-4}}{\pi(9.05\pi+\log m_{1})A_{4}(\alpha\log m_{2})\log^{2}(10545513\log m_{1})}.\end{split}$
@@
(139) $A_{4}(\alpha\log m_{2})\log^{2}(10545513\alpha\log m_{1})<0.785318\sqrt{m_{1}}$
@@
Now we show that $m_{2}<\exp(295490)$. Assume that $m_{2}\geq\exp(295490)$. If $m_{1}\geq M_{1}^{\prime}:=9.203\times 10^{22}$, then, since $m_{2}>e^{40000}$, we iterate the argument that we have just done with $(\tilde{m}_{1},\tilde{m}_{2})=(10^{14},e^{40000})$ to obtain (137) and (138) with $42.29385$, $116.73985$, and $441.0489$ in place of $150.9763$, $381.3734$, and $1922.630$. Eventually, we obtain
(140) $A_{4}(\alpha\log m_{2})\log^{2}(6665783\alpha\log m_{1})<0.785358\sqrt{m_{1}}$
with $A_{4}=1675.249$ instead of (139). On the other hand, if $m_{1}\leq 5\times 10^{14}G^{2}$, then, (120) gives $G<A_{4}(\log^{2}(5\times 10^{14}G^{2}))\log^{4}(6665783\log(2.56\times 10^{10}G^{2}))$ and therefore $G<3\times 10^{12}$, which implies that $m_{2}>e^{40000}>5\times 10^{14}G^{2}$. Hence, in any case, we must have $m_{2}>5\times 10^{14}G^{2}$ to obtain (123) and then (126) with $10^{-7}$ in place of $7.9\times 10^{-5}$ and $0.785398$ in place of $0.785319$, contrary to (140). Hence, we must have $m_{1}<M_{1}^{\prime}$ when $m_{2}\geq e^{40000}$.
@@
Hence, we must have $m_{1}<2.549\times 10^{24}$ and $m_{2}<\exp(295490)$. Moreover, if $m_{1}\geq 9.203\times 10^{22}$, then $m_{2}<e^{40000}$. If $m_{1}\geq 10^{14}$ and $m_{2}\geq e^{40000}$, then, Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(10^{14},e^{40000})$ yields that $\log x\leq(\log 2+k_{i}\log m_{1}+\ell_{i}\log m_{2})/2<5.07\times 10^{12}$, $x_{i}<\exp(5.07\times 10^{12})$, and $\left\lvert y_{i}\right\rvert\leq 2FG<1.316\times 10^{19}$. Similarly, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})$ according to Table 6 in each case, we can confirm that $x_{i}<\exp(5.07\times 10^{12})$ and $\left\lvert y_{i}\right\rvert<1.319\times 10^{19}$ whenever $m_{1}\geq 53$.
@@
$[10^{7},10^{14}]$ $\geq e^{40000}$ $(10^{7},e^{40000})$ $219940$ $2.992\times 10^{12}$ $1.010\times 10^{19}$
$[5000,10^{7}]$ $\geq e^{40000}$ $(5000,e^{40000})$ $161783$ $1.507\times 10^{12}$ $6.964\times 10^{18}$
$[53,5000]$ $\geq e^{40000}$ $(53,e^{40000})$ $139523$ $1.158\times 10^{12}$ $9.027\times 10^{18}$
$\geq 5000$ $[e^{2000},e^{40000}]$ $(5000,e^{2000})$ $40000$ $1.475\times 10^{12}$ $7.736\times 10^{18}$
@@
$\geq 53$ $[m_{1}^{*},e^{2000}]$ $(53,m_{1}^{*})$ $2000$ $1.830\times 10^{11}$ $2.890\times 10^{18}$
$41$ $\geq e^{40000}$ $(41,e^{40000})$ $118600$ $3.674\times 10^{11}$ $2.451\times 10^{18}$
$37$ $\geq e^{40000}$ $(37,e^{40000})$ $109518$ $1.346\times 10^{11}$ $3.663\times 10^{17}$
$29$ $\geq e^{40000}$ $(29,e^{40000})$ $113953$ $2.227\times 10^{11}$ $1.034\times 10^{18}$
$17$ $\geq e^{40000}$ $(17,e^{40000})$ $110578$ $1.501\times 10^{11}$ $5.745\times 10^{17}$
$13$ $\geq e^{40000}$ $(13,e^{40000})$ $116817$ $2.968\times 10^{11}$ $2.352\times 10^{18}$
$5$ $\geq e^{40000}$ $(5,e^{40000})$ $115906$ $2.180\times 10^{11}$ $2.038\times 10^{18}$

### v6

@@ 1. Introduction
    If $m_{1}m_{2}\geq 10^{10}$ and $m_{2}$ does not divide $x_{i}^{2}+1$ for some $i$, then $m_{1}<2.531\times 10^{24}$, $m_{2}<\exp(294622)$, $x_{i}<\exp(5.054\times 10^{12})$, and $\left\lvert y_{i}\right\rvert<1.312\times 10^{19}$. Moreover, $m_{2}<e^{40000}$ if additionally $m_{1}\geq 9.134\times 10^{22}$.
@@ 3. A lower bound for linear forms of three logarithms
(92) $\begin{split}{b^{\prime}}=&\frac{b_{1}}{9\pi+2D\operatorname*{h}(\alpha)}+\frac{b_{2}}{9\pi},\\ D=&[{\mathbb{Q}}(\alpha):{\mathbb{Q}}]/2,\\ a=&9\pi+2D\operatorname*{h}(\alpha),\\ H=&\max\{17,D,D(\log{b^{\prime}}+2.96)+0.01\}.\end{split}$
@@
(93) $\log\left\lvert\Lambda_{1}\right\rvert>-(2.76701a+0.12945)H^{2}.$
@@
This follows from Theorem 1 of [6] with $D=1$ by setting $\mu=0.59$, $\rho_{1}=18$, $\alpha_{1}=i$, $\alpha_{2}=\alpha$, $a_{1}=\rho\pi/2$, $a_{2}=2D\operatorname*{h}(\alpha)+(\rho\pi/2)$, $L=\left\lfloor H+\sqrt{H^{2}+(1/4)}+(1/2)\right\rfloor$, $k=(v_{1}(L)+\sqrt{v_{1}(L)^{2}+4v_{0}(L)v_{2}(L)})^{2}/(2v_{2}(L))^{2}$ to be the positive real number such that $\sqrt{k}$ satisfies the quadratic equation $v_{2}(L)k-v_{1}(L)\sqrt{k}-v_{0}(L)=0$, where
@@
Then, we can confirm that $0.27910.2437<\sqrt{k}<$ 0.2791, $H>7.5$, $15\leq L<0.92h$, and $K>712$ and then the condition (2) of Theorem 1 of [6]. Moreover, we can also confirm that $R_{2}-1<2a_{2}H^{2}<b_{1}$ and $S_{2}-1<2a_{1}H^{2}<b_{2}$. Hence the condition (1) of Theorem 1 of [6] also holds. Then, confirming that $\sqrt{k}L<0.239537H$, Theorem 1 of [6] yields that $\log\left\lvert\Lambda_{1}\max\{LSe^{LS\left\lvert\Lambda_{1}\right\rvert/(2b_{2})}/(2b_{2}),LRe^{LR\left\lvert\Lambda_{1}\right\rvert/(2b_{1})}/(2b_{1})\}\right\rvert>-(2.7666a+0.10984)H^{2}$ and then $\log\left\lvert\Lambda_{1}\right\rvert>-(2.76701a+0.12945)aH^{2}$. For the detail of the proof, see [22]. ∎
@@ 5. Proof of the Theorem
We begin by showing that $m_{1}<M_{1}$ with $M_{1}=2.531\times 10^{24}$. Assume that $m_{2}\geq m_{1}\geq M_{1}$. Now, Theorem 4.1 allows us to take $A_{4}=1752.029$, $A_{5}=10545512.8\cdots$, and $\rho=7.292$.
@@
Now, if $m_{1}\leq 7.2\times 10^{8}G^{2}$, then, (120) gives $m_{1}<7.2\times 10^{8}A_{4}^{2}(\log^{2}m_{1})\log^{4}(10545513\log m_{1})$ and therefore we must have $m_{1}<1.2\times 10^{24}<M_{1}$. Thus, we may assume that $m_{2}>m_{1}>7.2\times 10^{8}G^{2}$.
@@
(123) $\left\lvert\frac{r\pi}{4}-y_{1}\arctan\frac{1}{x_{1}}\right\rvert<\frac{k_{1}\ell_{3}}{x_{2}-1}+\frac{k_{1}\ell_{2}}{x_{3}-1}<\frac{(2+10^{-5})G}{\sqrt{m_{2}}}<7.5\times 10^{-5}.$
@@
(124) $\frac{\left\lvert y_{1}\right\rvert}{\sqrt{m_{1}}}>(1-10^{-9})\left(\frac{\pi}{4}-7.5\times 10^{-5}\right)>0.785323.$
@@
(129) $\log\left\lvert\Lambda_{1}\right\rvert>-(2.76701a+0.12945)aH^{2},$
@@
(130) $\begin{split}a=&9\pi+\log m_{1},\\ H=&\max\left\{17,2.97+\log\left(\frac{\left\lvert r\right\rvert}{9\pi+\log m_{1}}+\frac{\left\lvert y_{1}\right\rvert}{9\pi}\right)\right\}.\end{split}$
@@
(131) $\left\lvert\frac{r\pi}{4}\right\rvert-7.5\times 10^{-5}\leq\frac{\left\lvert y_{1}\right\rvert}{x_{1}-1}<\frac{A_{4}(\alpha\log m_{2})\log^{2}(10545513\log m_{1})}{m_{1}^{1/2}}.$
@@
(132) $\frac{\left\lvert r\right\rvert}{9\pi+\log m_{1}}+\frac{\left\lvert y_{1}\right\rvert}{9\pi}<A_{4}\psi_{3}(m_{1})(\alpha\log m_{2})\log^{2}(10545513\log m_{1}),$
@@
(133) $\begin{split}\psi_{3}(m_{1})=&\leavevmode\nobreak\ \frac{4}{(9\pi+\log m_{1})\pi m_{1}^{1/2}}+\frac{1}{9\pi}\\ &\leavevmode\nobreak\ +\frac{3.16\times 10^{-4}}{\pi(9\pi+\log m_{1})A_{4}(\alpha\log m_{2})\log^{2}(10545513\log m_{1})}.\end{split}$
@@
(139) $A_{4}(\alpha\log m_{2})\log^{2}(10545513\alpha\log m_{1})<0.785323\sqrt{m_{1}}$
@@
Now we show that $m_{2}<\exp(294622)$. Assume that $m_{2}\geq\exp(294622)$. If $m_{1}\geq M_{1}^{\prime}:=9.134\times 10^{22}$, then, since $m_{2}>e^{40000}$, we iterate the argument that we have just done with $(\tilde{m}_{1},\tilde{m}_{2})=(10^{14},e^{40000})$ to obtain (137) and (138) with $42.16707$, $116.41714$, and $439.5026$ in place of $150.5379$, $380.3347$, and $1916.227$. Eventually, we obtain
(140) $A_{4}(\alpha\log m_{2})\log^{2}(6665783\alpha\log m_{1})<0.785376\sqrt{m_{1}}$
with $A_{4}=1675.249$ instead of (139). On the other hand, if $m_{1}\leq 5\times 10^{14}G^{2}$, then, (120) gives $G<A_{4}(\log^{2}(5\times 10^{14}G^{2}))\log^{4}(6665783\log(2.56\times 10^{10}G^{2}))$ and therefore $G<3\times 10^{12}$, which implies that $m_{2}>e^{40000}>5\times 10^{14}G^{2}$. Hence, in any case, we must have $m_{2}>5\times 10^{14}G^{2}$ to obtain (123) and then (126) with $10^{-6}$ in place of $7.5\times 10^{-5}$ and $0.785376$ in place of $0.785323$, contrary to (140). Hence, we must have $m_{1}<M_{1}^{\prime}$ when $m_{2}\geq e^{40000}$.
@@
Hence, we must have $m_{1}<2.531\times 10^{24}$ and $m_{2}<\exp(294622)$. Moreover, if $m_{1}\geq 9.134\times 10^{22}$, then $m_{2}<e^{40000}$. If $m_{1}\geq 10^{14}$ and $m_{2}\geq e^{40000}$, then, Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})=(10^{14},e^{40000})$ yields that $\log x\leq(\log 2+k_{i}\log m_{1}+\ell_{i}\log m_{2})/2<5.07\times 10^{12}$, $x_{i}<\exp(5.07\times 10^{12})$, and $\left\lvert y_{i}\right\rvert\leq 2FG<1.316\times 10^{19}$. Similarly, applying Theorem 4.1 with $(\tilde{m}_{1},\tilde{m}_{2})$ according to Table 6 in each case, we can confirm that $x_{i}<\exp(5.07\times 10^{12})$ and $\left\lvert y_{i}\right\rvert<1.319\times 10^{19}$ whenever $m_{1}\geq 53$.
@@
$[10^{7},10^{14}]$ $\geq e^{40000}$ $(10^{7},e^{40000})$ $219207$ $2.982\times 10^{12}$ $1.007\times 10^{19}$
$[5000,10^{7}]$ $\geq e^{40000}$ $(5000,e^{40000})$ $161128$ $1.501\times 10^{12}$ $6.935\times 10^{18}$
$[53,5000]$ $\geq e^{40000}$ $(53,e^{40000})$ $138880$ $1.153\times 10^{12}$ $8.986\times 10^{18}$
$\geq 5000$ $[e^{2000},e^{40000}]$ $(5000,e^{2000})$ $40000$ $1.475\times 10^{12}$ $7.735\times 10^{18}$
@@
$\geq 53$ $[m_{1}^{*},e^{2000}]$ $(53,m_{1}^{*})$ $2000$ $1.830\times 10^{11}$ $2.889\times 10^{18}$
$41$ $\geq e^{40000}$ $(41,e^{40000})$ $118023$ $3.655\times 10^{11}$ $2.438\times 10^{18}$
$37$ $\geq e^{40000}$ $(37,e^{40000})$ $108986$ $1.339\times 10^{11}$ $3.644\times 10^{17}$
$29$ $\geq e^{40000}$ $(29,e^{40000})$ $113396$ $2.215\times 10^{11}$ $1.028\times 10^{18}$
$17$ $\geq e^{40000}$ $(17,e^{40000})$ $110031$ $1.493\times 10^{11}$ $5.715\times 10^{17}$
$13$ $\geq e^{40000}$ $(13,e^{40000})$ $116234$ $2.953\times 10^{11}$ $2.339\times 10^{18}$
$5$ $\geq e^{40000}$ $(5,e^{40000})$ $115315$ $2.169\times 10^{11}$ $2.027\times 10^{18}$